The scores of a college entrance examination had a normal distribution with mean u=552.7 and standard deviation of 28.6.

(a) What is the probability that a single student randomly chosen from all those who took the test had a score of 557 or higher?
Answer #1

X ~ N ( µ = 552.7 , σ = 28.6 )

We covert this to standard normal as

P ( X < x) = P ( (Z < X - µ ) / σ )

P ( X > 557 ) = P(Z > (557 - 552.7 ) / 28.6 )

= P ( Z > 0.15 )

= 1 - P ( Z < 0.15 )

= 1 - 0.5596 (From Z table)

= 0.4404